What weights of P4O6 and P4O10 will be produced by the combustion of 31g of P4 and 32g of oxygen leaving no P4 and O2

Grade:11

2 Answers

Arun
4 years ago
P4O6 = 56.3437% P 
P4O10 = 43.6420% P 
31 g P = 31 / (31 + 32) = 49.2063% P 
So it becomes a problem in proportions: 
100% P4O6 = 56.3437% 
z% P4O6 = 49.2063% 
0% P4O6 = 43.6420% 
Find z. 
z / 100 = (49.2063 - 43.6420) / (56.3437 - 43.6420) 
Solve for z algebraically: 
z = 43.808% P4O6 
By the Law of the Conservation of Matter, the total mass 
of P4O6 and P4O10 = 31g + 32g = 63g 
43.808% of 63g = 27.6 g P4O6 
63g - 27.6g = 35.4 g P4O10
 
 
Khimraj
4 years ago
It is based on limiting reagent
here 2P4 + 8O2 gives P4O6+ P4O10
so phosphorous having 31g and oxygen having 32g
so P4 is the limiting reagent
so ignoring oxygen 
2P4----- gives P4O6
so 2x124 gives 1220
so 31 gives 152.5
so the answer is 152.5